OpenStudy (anonymous):

Part 1: Decide whether you would Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3 Can someone help me? Thanks!

OpenStudy (anonymous):

Why don't you pick one and we'll try it.

OpenStudy (anonymous):

OK, let's do a graph.

OpenStudy (anonymous):

substitution

OpenStudy (anonymous):

I got (2,-3) as the place where the two lines intersect. Plug those into the equations to check my work. Looks good. Are you still here @CaseyCarns

OpenStudy (anonymous):

since 3x+ y =3 can be y = -3x+3 so 4x + 3(-3x+3) = -1

OpenStudy (anonymous):

Yeah sorry

OpenStudy (anonymous):

Substitution. y = 3 - 3x from the second equation. Plug into the first: 4x + 3 (3 - 3x) = -1 4x + 9 - 9x = -1 -5x = -10 x = 2

OpenStudy (anonymous):

Do you understand the substitution method from the replies above or shall we work through it.

OpenStudy (anonymous):

and for elimination multiply 2nd eq, by 3 to eliminate y

OpenStudy (anonymous):

I do thank you so much!

OpenStudy (anonymous):

All good then. yw